# Flat band mediated photon-photon interactions in 2D waveguide QED networks

### v2

@@ III.2 Compact localized states
where we omitted the band index. If there exists a set of coefficients $\alpha_{\mathbf{k}}$ such that $A^{(\mathbf{R})}_{{\mathbf{R}^{\prime},\beta}}$ is nonzero only for a finite number of unit cells around $\mathbf{R}$, then $|\chi^{(\mathbf{R})}\rangle$ is called a compact localized state. This implies that each component $\alpha_{\mathbf{k}}v_{\mathbf{k},\beta}$ must be given by a finite sum of Bloch phases. Sufficient conditions to fulfill this requirement are the presence of a flat band and finite-range hoppings in the Hamiltonian [9, 10, 8]. Importantly, while the latter condition is not met in our setting as it hosts long-range hoppings, it is still possible to find suitable coefficients $\alpha_{\mathbf{k}}$ such that $\alpha_{\mathbf{k}}v_{\mathbf{k},\beta}$ becomes a finite sum of Bloch phases (see App. A). This is possible again due to destructive interference of the plan-waves-mediated interactions in the network. A generic compact localized state centered around the lattice position $\mathbf{R_{0}}$ is given by:
@@ III.3 Quantum geometric tensor
$\begin{split}T_{xx}(k_{x},k_{y})&=\frac{\big{|}t_{y}\big{|}^{2}\big{|}\partial_{k_{x}}t_{x}\big{|}^{2}}{\Big{(}|t_{x}|^{2}+|t_{y}|^{2}\Big{)}^{2}}\in\mathbb{R},\\ T_{yy}(k_{x},k_{y})&=\frac{\big{|}t_{x}\big{|}^{2}\big{|}\partial_{k_{y}}t_{y}\big{|}^{2}}{\Big{(}|t_{x}|^{2}+|t_{y}|^{2}\Big{)}^{2}}\in\mathbb{R},\\ T_{xy}(k_{x},k_{y})&=-\frac{t_{x}^{*}(\partial_{k_{x}}t_{x})\,t_{y}(\partial_{k_{y}}t_{y})^{*}}{\Big{(}|t_{x}|^{2}+|t_{y}|^{2}\Big{)}^{2}}\,,\end{split}$ (13)
@@ IV Interaction induced transport
Figure 3: The two-excitation dispersion in the softcore interactions regime within the isolated-band approximation calculated by diagonalizing the matrix given in Eq. (16) as a function of the center-of-mass momentum $\mathbf{K}$. Panel (a) displays half of the full 2D Brillouin zone ($k_{y}<0$) for better resolution, while panel (b) shows the dispersion along the symmetry lines of the Brillouin zone. (c) Relative coordinate population distribution of the bound states having the interaction-induced dispersive bands shown in panles (a)-(b) at the $M,\Gamma$ and $X$ symmetry-points. The upper (lower) panel displays the upper (lower) branch of the bound states. The color map represents the probability distribution of finding the second excitation on different atoms in the lattice, given that the first excitation is located on the atom marked with the orange circle. The results were obtained for a system of $30\times 30$ unit cells.
@@ IV.1.1 Two excitation spectrum
We start by examining the softcore interaction regime, where the band separation is much larger than the emitter nonlinearity that governs effective photon-photon interactions, $U\ll\Delta\epsilon$. This allows us to use the isolated-band approximation [22], which assumes that the interaction is sufficiently weak to prevent coupling between states from different bands, enabling us to project the two-excitation Hamiltonian onto the flat-band states alone. This can be done by expressing the non-interacting two-excitation flat-band states in terms of the center-of-mass (CM) coordinate, $\mathbf{R}=\frac{\mathbf{R}_{1}+\mathbf{R}_{2}}{2}$, and the relative coordinate, $\mathbf{r}=\mathbf{R}_{1}-\mathbf{R}_{2}$. We thus obtain:
$\begin{split}&|\psi_{FB}^{(2)}(\mathbf{K},\mathbf{q})\rangle=\mathcal{N}^{(2)}\sum_{\mathbf{R,r}\atop\beta_{1}\beta_{2}}v_{\frac{\mathbf{K}}{2}+\mathbf{q},\beta_{1}}v_{\frac{\mathbf{K}}{2}-\mathbf{q},\beta_{2}}\times\\ &e^{i\left(\mathbf{K}\cdot\mathbf{R}+\mathbf{q}\cdot\mathbf{r}\right)}|\mathbf{R}-\frac{r}{2}\rangle|\beta_{1}\rangle\bigotimes|\mathbf{R}+\frac{r}{2}\rangle|\beta_{2}\rangle,\\ \end{split}$ (15)
where $\mathcal{N}^{(2)}$ is the states’ norm while $\mathbf{K}=\mathbf{k_{1}}+\mathbf{k_{2}}$ and $\mathbf{q}=\frac{k_{1}-k_{2}}{2}$ respectively the center of mass and relative momentum. Since the interaction term in Eq. (1) is translationally invariant, the CM momentum $\mathbf{K}$ is a good quantum number and the two-excitation Hamiltonian projected onto the flat-band states (15) can be diagonalized for each value of $\mathbf{K}$ independently:
$\begin{split}\hat{V}_{\mathbf{q},\mathbf{q^{\prime}}}(\mathbf{K})&=\langle\psi_{FB}(\mathbf{K},\mathbf{q^{\prime}})|\hat{H}_{int}|\psi_{FB}(\mathbf{K,q})\rangle\\ &=\sum_{\beta}v_{-\mathbf{\frac{K}{2}-q^{\prime}},\beta}v_{-\mathbf{\frac{K}{2}+q^{\prime}},\beta}v_{\mathbf{\frac{K}{2}+q},\beta}v_{\mathbf{\frac{K}{2}-q},\beta}.\end{split}$ (16)
In Fig. 3, we present the dispersion obtained by the diagonalization of Eq. (16) as a function of the center-of-mass momentum $\mathbf{K}$. The spectrum reveals $N-2$ degenerate zero-energy states for each $\mathbf{K}$, along with two dispersive branches that merge along the entire Brillouin zone boundary. The presence of these dispersive branches underscores the interaction-induced mobility of initially localized flat-band states. These states correspond to bound photon-photon pairs, as they appear within the energy gap formed by the degenerate flat band. Notice that these bound states obtained via the projected Hamiltonian are actually immersed in a continuum of states of the full Hamiltonian. Such a continuum originates from the the upper/lower single-particle polariton branches where each contributes to a single free photon to the two-excitation sector. We will show below that the bound states are however well-decoupled from such a continuum of modes.
@@ IV.1.2 Dynamics
Here the operator $\hat{c}_{\mathbf{R}_{0}}^{\dagger}$ generates a CLS centered at lattice site $\mathbf{R}_{0}$ as defined in Eq. (11) and $|0\rangle$ is the vacuum state. In the non-interacting regime ($U=0$) the state (17) is an eigenstate of the system and thus remains unchanged during the evolution. To study the dynamics of this state in the presence of interactions, we evolve it under the Schrödinger equation governed by the effective Hamiltonian (1) and projected onto the two-excitation subspace:
@@
Figure 4: Time evolution of the state defined in Eq. (17) in the softcore interactions regime with $U=0.1\gamma$. (a) Initial state fidelity $\mathcal{F}_{0}(t)$ (blue solid line), flat-band subspace projection $\mathcal{P}_{FB}$ (dashed red line) and norm of the evolved state $\mathcal{N}(t)$ (black dotted line) as a function of time. (b) Snapshots of the excitation population evolution at times (a) $t=0$, (b) $t=t_{min}\left[\mathcal{F}_{0}\right]$ and (c) $\gamma t=10^{4}$, for two initially excited neighboring CLSs. The simulation was performed for a system size of $8\times 8$ unit cells. (c) Projection of the two-excitation state onto the interaction-induced dispersive bound states (red solid line) and onto the interacting non-dispersive states (gray dashed line) shown in Fig. 3.
@@
In Ref. [22] a connection between isolated flat bands and the inverse mass tensor ($m_{ij}^{-1}$) of interacting low-energy bound states was shown to involve the quantum metric and, in particular, its integral via $m_{ij}^{-1}\sim\int\text{d}\mathbf{k}\,\text{Re}(T_{ij}(\mathbf{k}))$, which in our case we found to be finite (see Eq. (14)). Whether this connection is also occurring here, and under what conditions, remains however an open question. In particular, we would like to highlight that Ref. [22] established this connection only for a single bound-state branch, despite multiple branches, each separated by an energy gap, occurred in the two-body spectrum. In our case, the multiple bound-state branches are energy degenerate, see Fig. 3(b), which implicitly introduces the notion of an emergent sublattice for the bound pairs. A similar phenomenon was recently shown to occur at high energy for strongly-interacting nearest-neighbor bound pairs [113] based on the concept of line-graph lattice for their center of mass. This provided a clear geometrical explanation behind the origin of multi-band bound states branches. Here in the softcore interactions regime a less straightforward mechanism might be in place. This fact prevents a plain-vanilla employment of the reasonings discussed in Ref. [22] to connect the bound states mass to the quantum metric, and possibly requiring a non-abelian/multi-branch treatment.
Figure 5: (a) Same as in Fig. 4(a) in the hardcore interactions regime ($U\rightarrow\infty)$, for the initial condition defined in Eq. (21). (b) Scaling of the fidelity oscillation frequency, $\omega_{0}$ for the initial condition defined in Eq. (21), as a function of interaction strength $U$.
@@ IV.2 Two level atom limit
$\hat{s}^{\dagger}_{\mathbf{R_{0}}}=\frac{1}{2}\Big{(}\hat{\sigma}^{+}_{\mathbf{R_{0}},A}+\hat{\sigma}^{+}_{\mathbf{R_{0}}-d\mathbf{\hat{y},A}}-\hat{\sigma}^{+}_{\mathbf{R_{0}},C}-\hat{\sigma}^{+}_{\mathbf{R_{0}}-d\mathbf{\hat{x},C}}\Big{)}.$ (22)
@@ V Conclusion
In this paper, we have shown that in a waveguide QED system consisting of a 2D Lieb-like array of quantum emitters coupled to a network of 1D waveguides, infinitely long-range interactions can give rise to an energetically isolated flat band. We prove that this band is nontrivial, possessing nonvanishing geometric tensor components and hosting perfectly subradiant compact localized states. We then demonstrate that for two excitations, emitter nonlinearity can induce strong effective photon-photon interactions, leading to interaction-induced transport of bound photon pairs. In the softcore interactions regime, propagation occurs at a rate that scales linearly with the effective interaction strength, mirroring predictions for fermionic systems [22]. This raises the intriguing question whether in the half-filling regime [7] this platform could host superfluid states [21, 8, 9, 10]. The ability to couple superconducting qubits to transmission lines in various architectures [42, 43, 44, 45, 46] makes the microwave regime a promising platform for experimentally probing our findings. Additionally, the tunability of the emitters nonlinearity would enable the exploration of the different interacting regimes discussed in this work. Alternatively, our results could be probed in the optical regime by interfacing waveguide networks with cold or artificial atoms [36, 37, 38, 39, 40], or, in the single-excitation linear regime, by utilizing other photonic modes  [114]. An intriguing open questions is about the impact of long-range couplings either in topological bands or in generating purely interacting topological bound states [113].
@@ Appendix A Compact localized states derivation
$\begin{split}1+i\tan\left(\frac{k_{i}d}{2}\right)&=\frac{\cos\left(\frac{k_{i}d}{2}\right)+i\sin\left(\frac{k_{i}d}{2}\right)}{\cos{\left(\frac{k_{i}d}{2}\right)}}\\ &=2\frac{e^{i\frac{k_{i}d}{2}}}{e^{i\frac{k_{i}d}{2}}+e^{i\frac{k_{i}d}{2}}}=\frac{2}{1+e^{-ik_{i}d}}\,,\end{split}$ (24)

### v3

@@ III.2 Compact localized states
where we omitted the band index. If there exists a set of coefficients $\alpha_{\mathbf{k}}$ such that $A^{(\mathbf{R})}_{{\mathbf{R}^{\prime},\beta}}$ is nonzero only for a finite number of unit cells around $\mathbf{R}$, then $|\chi^{(\mathbf{R})}\rangle$ is called a compact localized state. This implies that each component $\alpha_{\mathbf{k}}v_{\mathbf{k},\beta}$ must be given by a finite sum of Bloch phases. Sufficient conditions to fulfill this requirement are the presence of a flat band and finite-range hoppings in the Hamiltonian [9, 10, 8]. Importantly, while the latter condition is not met in our setting as it hosts long-range hoppings, it is still possible to find suitable coefficients $\alpha_{\mathbf{k}}$ such that $\alpha_{\mathbf{k}}v_{\mathbf{k},\beta}$ becomes a finite sum of Bloch phases (see App. A). This is possible again due to destructive interference of the plane-waves-mediated interactions in the network. A generic compact localized state centered around the lattice position $\mathbf{R_{0}}$ is given by:
@@ III.3 Quantum geometric tensor
[…]
@@ IV Interaction induced transport
Figure 3: The two-excitation dispersion in the softcore interactions regime within the isolated-band approximation calculated by diagonalizing the matrix given in Eq. (23) as a function of the center-of-mass momentum $\mathbf{K}$. Panel (a) displays half of the full 2D Brillouin zone ($k_{y}<0$) for better resolution, while panel (b) shows the dispersion along the symmetry lines of the Brillouin zone. (c) Relative coordinate population distribution of the bound states having the interaction-induced dispersive bands shown in panls (a)-(b) at the $M,\Gamma$ and $X$ symmetry-points. The upper (lower) panel displays the upper (lower) branch of the bound states. The color map represents the probability distribution of finding the second excitation on different atoms in the lattice, given that the first excitation is located on the atom marked with the orange circle. The results were obtained for a system of $30\times 30$ unit cells.
@@ IV.1.1 Two excitation spectrum
We start by examining the softcore interaction regime, where the band separation is much larger than the emitter nonlinearity that governs effective photon-photon interactions, $U\ll\Delta\epsilon$. This allows us to use the isolated-band approximation [23], which assumes that the interaction is sufficiently weak to prevent coupling between states from different bands, enabling us to project the two-excitation Hamiltonian onto the flat-band states alone. Higher-excitation flat-band states can be constructed as product states of non-overlapping CLSs, which are exact zero-energy eigenstates of $\hat{H}_{\mathrm{eff}}$, independent of the value of $U$. Specifically, the non-interacting two-excitation flat-band states can be expressed in terms of the center-of-mass (CM) coordinate, $\mathbf{R}=\frac{\mathbf{R}_{1}+\mathbf{R}_{2}}{2}$, and the relative coordinate, $\mathbf{r}=\mathbf{R}_{1}-\mathbf{R}_{2}$. We thus obtain:
[…]
where $\mathcal{N}^{(2)}$ is the states’ norm while $\mathbf{K}=\mathbf{k_{1}}+\mathbf{k_{2}}$ and $\mathbf{q}=\frac{k_{1}-k_{2}}{2}$ are, respectively, the center of mass momentum and relative momentum. To account for the effect of the interaction, we note that the interaction term in Eq. (9) is translationally invariant, and therefore the center-of-mass momentum $\mathbf{K}$ is a good quantum number. The two-excitation Hamiltonian projected onto the flat-band states (22) can then be diagonalized independently for each value of $\mathbf{K}$:
$\begin{split}\hat{V}_{\mathbf{q},\mathbf{q^{\prime}}}(\mathbf{K})&=\langle\psi_{FB}(\mathbf{K},\mathbf{q^{\prime}})|\hat{H}_{int}|\psi_{FB}(\mathbf{K,q})\rangle\\ &=U\sum_{\beta}v_{-\mathbf{\frac{K}{2}-q^{\prime}},\beta}v_{-\mathbf{\frac{K}{2}+q^{\prime}},\beta}v_{\mathbf{\frac{K}{2}+q},\beta}v_{\mathbf{\frac{K}{2}-q},\beta}.\end{split}$ (23)
In Fig. 3, we present the dispersion $\epsilon^{(2)}(\mathbf{K})$ obtained by the diagonalization of Eq. (23) as a function of the center-of-mass momentum $\mathbf{K}$. The spectrum reveals N2N-2 degenerate zero-energy states for each $\mathbf{K}$, along with two dispersive branches that merge along the entire Brillouin zone boundary. The presence of these dispersive branches underscores the interaction-induced mobility of initially localized flat-band states. These states correspond to bound photon-photon pairs, as they appear within the energy gap formed by the degenerate flat band. Notice that these bound states obtained via the projected Hamiltonian are actually immersed in a continuum of states of the full Hamiltonian. Such a continuum originates from the upper and lower single-particle polariton branches where each contributes to a single free photon to the two-excitation sector. We will show below that the bound states are however well-decoupled from such a continuum of modes.
@@ IV.1.2 Dynamics
Here the operator $\hat{c}_{\mathbf{R}_{0}}^{\dagger}$ generates a CLS centered at lattice site $\mathbf{R}_{0}$ as defined in Eq. (18) and $|0\rangle$ is the vacuum state. In the non-interacting regime ($U=0$) the state (24) is an eigenstate of the system and thus remains unchanged during the evolution. To study the dynamics of this state in the presence of interactions, we consider the effective Hamiltonian (9) projected onto the two-excitation subspace, $\hat{H}^{(2)}_{\text{eff}}$, and we evolve the state under the Schrödinger equation:
@@
Figure 4: Time evolution of the state defined in Eq. (24) in the softcore interactions regime with $U=0.1\gamma$. (a) Initial state fidelity $\mathcal{F}_{0}(t)$ (blue solid line), flat-band subspace projection $\mathcal{P}_{FB}$ (dashed red line) and norm of the evolved state $\mathcal{N}(t)$ (black dotted line) as a function of time. (b) Snapshots of the excitation population evolution at times $t=0$ (left panel), $t=t_{min}\left[\mathcal{F}_{0}\right]$ (center panel) and $\gamma t=10^{4}$ (right panel), for two initially excited neighboring CLSs. The simulation was performed for a system size of $8\times 8$ unit cells. (c) Projection of the two-excitation state onto the interaction-induced dispersive bound states (red solid line) and onto the interacting non-dispersive states (gray dashed line) shown in Fig. 3.
@@
In Ref. [23] a connection between isolated flat bands and the inverse mass tensor ($m_{ij}^{-1}$) of interacting low-energy bound states was shown to involve the quantum metric and, in particular, its integral via $m_{ij}^{-1}\sim\int\text{d}\mathbf{k}\,\text{Re}(T_{ij}(\mathbf{k}))$, which in our case we found to be finite (see Eq. (21)). Whether this connection is also occurring here, and under what conditions, remains however an open question. In particular, we would like to highlight that Ref. [23] established this connection only for a single bound-state branch, despite the fact that multiple branches, each separated by an energy gap, occurred in the two-body spectrum. In our case, the multiple bound-state branches are energy intertwined, see Fig. 3(b), which implicitly introduces the notion of an emergent flavor for the bound pairs. A similar phenomenon was recently shown to occur at high energy for strongly-interacting nearest-neighbor bound pairs [111] based on the concept of line-graph lattice for their center of mass. This provided a clear geometrical explanation behind the origin of multi-band bound states branches. Here in the softcore interactions regime a less straightforward mechanism might be in place originating from flat band mediated correlations. This fact prevents a plain-vanilla employment of the reasonings discussed in Ref. [23] to connect the bound states mass to the quantum metric, and possibly requiring a non-abelian/multi-branch treatment to include the bound state flavors into the treatment.
Figure 5: (a) Same as in Fig. 4(a) in the hardcore interactions regime ($U\rightarrow\infty)$, for the initial condition defined in Eq. (28). (b) Scaling of the fidelity oscillation frequency, $\omega_{0}$ for the initial condition defined in Eq. (28), as a function of interaction strength $U$. The continuous line indicates the fit of the numerical data in the linear regime.
@@ IV.2 Two level atom limit
[…]
@@ V Conclusion
In this paper, we have shown that in a waveguide QED system consisting of a 2D Lieb-like array of quantum emitters coupled to a network of 1D waveguides, infinitely long-range interactions can give rise to an energetically isolated flat band. We prove that this band is nontrivial, possessing nonvanishing geometric tensor components and hosting perfectly subradiant compact localized states. We then demonstrate that for two excitations, emitter nonlinearity can induce strong effective photon-photon interactions, leading to interaction-induced transport of bound photon pairs. In the softcore interactions regime, propagation occurs at a rate that scales linearly with the effective interaction strength, mirroring predictions for fermionic systems [23]. This raises the intriguing question whether in the half-filling regime [7] this platform could host superfluid states [22, 8, 9, 10]. The ability to couple superconducting qubits to transmission lines in various architectures with negligible coupling to other dissipation channels [43, 44, 45, 46, 47] makes the microwave regime a promising platform for experimentally probing our findings. Additionally, the tunability of the emitters nonlinearity would enable the exploration of the different interacting regimes discussed in this work. The main limitation in this setting arises from the possible breakdown of the Markov approximation, which imposes an upper bound on the atom-field coupling, the lattice spacing and the number of emitters [112]. The former can be experimentally controlled by ensuring that the coupling frequency remains smaller than the bandwidth of the transmission line. The latter can be controlled by fixing the emitter spacing to $\lambda/2$, as required in our proposal. For experimental parameters similar to those in Ref. [47], this approximation is expected to hold except for very large arrays on the order of several hundred emitters. Alternatively, our results could be probed in the optical regime by interfacing waveguide networks with cold or artificial atoms [37, 38, 39, 40, 41], or, in the single-excitation linear regime, by utilizing other photonic modes [113]. In this scenario, the main challenges concern trapping atoms near complex photonic structures and the presence of decay into radiative modes other than the waveguide.
Finally, an intriguing open question is about the impact of long-range couplings either in topological bands or in generating purely interacting topological bound states [111].
@@ Appendix A Compact localized states derivation
[…]